Wilde Corporation budgeted the following costs for the production of its one and only product for the next fiscal​ year:

Direct materials ​$1,140,000

Direct labor ​795,000

Manufacturing overhead: Variable ​840,000   // Fixed ​700,000

Selling and administrative: Variable ​360,000 // Fixed ​530,000

Total costs ​$4,365,000

Wilde has an annual target operating income of​ $920,000.

The markup percentage for setting prices as a percentage of the variable cost of the product is​ ________.

A. 68.6% [right answer]

B. 27.5%

C. 39.2%

D. 46.3%

I know the correct answer, but not how to arrive at that answer. Please explain. Thank you!
Answer #1

Total fixed cost = 700,000+530,000 = 1,230,000

Total Variable cost = 1,140,000+795,000+840,000+360,000 = 3,135,000

The markup percentage for setting prices as a percentage of the variable cost of the product

= (Total fixed cost + Target operating income) /Total Variable cost

= (1,230,000+920,000)/3,135,000

= 2,150,000/3,135,000

= 68.6%